OpenStudy (anonymous):

Solve for X. X/4 = 2 1/4 *

OpenStudy (anonymous):

Solve for X. 1/2 * X = 9 *

OpenStudy (igreen):

For the first one, multiply 4 to both sides.

OpenStudy (igreen):

\(\sf \dfrac{x}{4} = 2 \dfrac{1}{4}\) \(\sf \dfrac{x}{4} \times 4 = 2 \dfrac{1}{4} \times 4\) \(\sf \dfrac{x}{\cancel{4}} \times \cancel{4} = 2 \dfrac{1}{4} \times 4\) \(\sf x = 2\dfrac{1}{4} \times 4\) \(\sf x = \dfrac{9}{4} \times 4\) \(\sf x = \dfrac{9 \times 4}{4}\) Simplify that.

OpenStudy (igreen):

That's incorrect try again.

OpenStudy (cmtboy2):

no it's not

OpenStudy (cmtboy2):

no 2.1

OpenStudy (igreen):

No, that's incorrect.

OpenStudy (igreen):

Simplify \(\sf \dfrac{9 \times 4}{4}\)

OpenStudy (cmtboy2):

ya i got 2.1

OpenStudy (cmtboy2):

k12.1 it's 2.1/4